Sometimes you will want to include text exactly as it appears on a terminal screen. For example, you might want to include part of a computer program. Not only do you want L^{A}TEX to stop playing around with the layout of your text, you also want to be able to type all the characters on your keyboard without confusing L^{A}TEX. The verbatim environment has this effect:

     The section of program in
     question is:
     \begin{verbatim}
     { this finds %a & %b }
     for i := 1 to 27 do
       begin
       table[i] := fn(i);
       process(i)
       end;
     \end{verbatim}

The section of program in question is:

                                                     { this finds %a & %b }
                                                     for i := 1 to 27 do
                                                        begin
                                                        table[i] := fn(i);
                                                        process(i)
                                                        end;